OpenStudy (anonymous):

5(x – 5) = 4(x – 1) + 3x.

OpenStudy (anonymous):

distribute first

OpenStudy (anonymous):

5x-25=4x-4+3x

OpenStudy (anonymous):

5x=25 = 7x-4

OpenStudy (anonymous):

5x-25=7x +4

OpenStudy (anonymous):

5x-21=7x

OpenStudy (anonymous):

-21=7x -5x

OpenStudy (anonymous):

-21 = 2x / 2x

OpenStudy (anonymous):

-21/2

OpenStudy (anonymous):

thank Yuh